# Enhancing Retrosynthesis with Conformer: A Template-Free Method

### v1

@@ 4.3 Training Strategy
Data Augmentation
When molecules are converted into SMILES, different choices of root atoms can result in different SMILES representations, and this non-uniqueness may impede training of seq2seq models. Apart from unifying SMILES into a canonical form, another approach is to transform and enlarge the original SMILES sets through data augmentation. Drawing from some effective data augmentation tricks in prior research Dai et al. (2019); Tetko et al. (2020); Zhong et al. (2022), we explore two types of data augmentation approaches: Randomize SMILES and Root-align SMILES.
  • Randomize SMILES: By shuffling the order of atoms in a molecule or altering the priority of side chains, different SMILES representations can be generated. There can be up to n! (where n is the number of atoms in the molecule) possible SMILES to represent a molecule.
  • Root-align SMILES (R- SMILES): First, a root atom is randomly selected from product, and then product R-SMILES is generated based on this root atom. Next, the same root atom in reactants is identified based on atom mapping or substructure matching between product and reactants and generate reactants R-SMILES. R-SMILES effectively reduces the editing distance between product and reactants, ensuring a high sequence similarity and simplifying the prediction problem.
    R-product: Brc1ccc(C=O)cn1
    R-reactants: Brc1ccc(Br)cn1.C(N(C)C)=O
In contrast to other methods that directly expand dataset, we employ an on-the-fly approach to progressively augment data during training iterations. The specific procedure involves transforming the input canonical SMILES into Randomize SMILES (or Root-align SMILES) with a 50% probability in each training iteration. As iterations increase, the model gradually gains a broader perspective on permuted SMILES. This dynamic augmentation approach allows the model to focus more on the standard form (canonical SMILES) while regularizing with the permuted form (permuted SMILES), thereby improving the accuracy, stability, and generalization of model.
@@
RetroSim 52.9 73.8 81.2 88.1 37.3 54.7 63.3 74.1
GLN 64.2 79.1 85.2 90.0 52.5 69.0 75.6 83.7
LocalRetro 63.9 86.8 92.4 96.3 53.4 77.5 85.9 92.4
@@
RetroXpert 62.1 75.8 78.5 80.9 50.4 61.1 62.3 63.4
G2G 61.0 81.3 86.0 88.7 48.9 67.6 72.5 75.5
GraphRetro 63.9 81.5 85.2 88.1 53.7 68.3 72.2 75.5
RetroPrime 64.8 81.6 85.0 86.9 51.4 70.8 74.0 76.1
MEGAN 60.7 82.0 87.5 91.6 48.1 70.7 78.4 86.1
Graph2Edits 67.1 87.5 91.5 93.6 55.1 77.3 83.4 89.4
@@
SCROP 59.0 74.8 78.1 81.1 43.7 60.0 65.2 68.7
GET 57.4 71.3 74.8 77.4 44.9 58.8 62.4 65.9
Aug. Transformer - - - - 48.3 - 73.4 77.4
GTA - - - - 51.1 67.6 74.8 81.6
Graph2SMILES - - - - 52.9 66.5 70.0 72.9
Retroformer 64.0 82.5 86.7 90.2 53.2 71.1 76.6 82.1
our method (Ours) 65.3 84.0 88.3 91.6 53.6 73.7 80.2 86.1
our method (Ours, +R) 67.2 87.0 91.4 94.4 55.5 76.3 82.8 88.8
Table 1: Top-k accuracy for Retrosynthesis prediction on USPTO-50K. +R indicates R-SMILES augmentation. The metric values of other models in the table are taken from their original papers. Best performance is in bold.
     Model Top-k validity (%)
1 3 5 10
Graph2SMILES 99.4 90.9 84.9 74.9
RetroPrime 98.9 98.2 97.1 92.5
Retroformer 99.2 98.5 97.4 96.7
our method (Ours) 99.8 99.6 99.1 97.0
our method (Ours, +R) 99.8 99.4 98.6 95.4
Table 2: Top-k SMILES validity for Retrosynthesis prediction on USPTO-50K with reaction class unknown.
@@ 4.4 Overall Loss
The overall loss is composed of the language modeling R-Drop lossWu et al. (2021) $\mathcal{L}_{LM}$ for Retrosynthesis, along with the SMILES Alignment guidance loss $\mathcal{L}_{SA}$. R-drop Loss builds upon cross-entropy loss enhancing prediction consistency and model robustness by minimizing bidirectional KL-divergence of output distributions of any pair of sub-models sampled from dropout in model training.
@@
$\mathcal{L}=\mathcal{L}_{LM}^{(CE)}+\alpha\mathcal{L}_{LM}^{(KL)}+\mathcal{L}_{SA}$ (10)
where $\mathcal{L}_{LM}^{(CE)}$ is Cross-Entropy loss to minimize prediction error against actual outputs, $\mathcal{L}_{LM}^{(KL)}$ is KL-divergence in R-Drop loss to strive for outputs of models with varying Dropout to be as similar as possible, and we set $\alpha$ to 0.5. $\mathcal{L}_{SA}$ is Cross-Entropy loss between cross-attention score of last decoder layer and SMILES Alignment Map that we mention in Section 4.3.
@@ 5 Experiments
We use the popular Retrosynthesis benchmark dataset USPTO-50K Schneider et al. (2016), which contains 50,016 atom-mapped reactions, to evaluate our model. The same data split as Coley et al. (2017) is applied to our experiments, resulting in 40,008, 5,001, and 5,007 reactions for the training, validation, and test sets, respectively. Conformers are generated using RDKit with the stochastic optimization algorithm Merck Molecular Force Field (MMFF). We then apply the same algorithm in Wan et al. (2022) to extract the ground truth SMILES Alignment Map.
Setting
@@
Our model is built on vanilla Transformer, consisting of 8 encoder layers and 8 decoder layers with 8 attention heads. The model dimension is set to 512. We employ the Adam optimizer Kingma and Ba (2014) with $(\beta_{1},\beta_{2})=(0.9,0.98)$ and train the model with batch size of 16. During training, we set an early-stop of 7 epochs under a maximum of 1000 epochs to obtain the 7 best models and average their parameters. Finally, the whole training process took approximately 30 hours on a single NVIDIA GeForce RTX 4090 GPU.
Evaluation
@@
During inference, we use beam search Tillmann and Ney (2003) decode strategy with a beam size is 10. After that, we evaluate our model with Top-k accuracy, which considers it correct only if Top-k predicted reactants of the beam search results are the same as original test set. For molecule validity, we verify Top-k validity based on whether RDKit can recognize the Top-k predicted reactants.
Baseline
@@
We compare our proposed method, our method, with several strong baseline models from three types of Retrosynthesis method. We take RetroSim Coley et al. (2017), GLN Dai et al. (2019) and LocalRetro Chen and Jung (2021) to represent template-based methods. We take RetroXpert Yan et al. (2020), G2Gs Shi et al. (2020), GraphRetro Somnath et al. (2020), RetroPrime Wang et al. (2021), MEGAN Nijssen and Kok (2004) and Graph2Edits Zhong et al. (2023) to represent semi-template method. We take SCROP Zheng et al. (2019), Aug.Transformer Tetko et al. (2020), Graph2SMILES Zhong et al. (2023), GET Mao et al. (2021), GTA Seo et al. (2021) and Retroformer Wan et al. (2022) to represent template-free method. We did not choose pre-training methods for comparison as they are trained on larger datasets. We experimented with two training strategies for our proposed model: our method represents the model with Randomize SMILES data augmentation; our method+R represents the model with Root-align SMILES data augmentation.
@@
Figure 2: Case study for different structures. Each subfigure shows predictions of different methods: our method (Top), Transformer (Bottom).
5.1 Performance
Top-k Accuracy
As shown in Table 1, for the model with random augmentation our method, when reaction class is known, Top-1/Top-10 accuracy is 65.3%/91.6%, exceeding previous state-of-the-art template-free methods. When reaction class is unknown, our model achieves Top-1/Top-10 accuracy at 53.6%/86.1% and also outperforms existing template-free methods. It’s worth noting that model demonstrates competitive performance in both template-based and semi-template methods, even though they rely on predefined templates or RDKit editing. For the model with R-SMILES augmentation our method+R, it is astonishingly found that by further aligning the root of product and reactants, our model not only greatly surpasses state-of-art template-free methods, but also essentially exceeds other template-based methods and semi-template methods on several metrics.
Top-k Validity
We take three typical template-free methods Graph2SMILES, RetroPrime and Retroformer as baseline models to verify the predicted reactants SMILES validity of our model. We refrained from selecting template-based and semi-template methods, as SMILES constructed based on template information ensure validity. Conversely, template-free methods often struggle to generate valid SMILES without extra chemical guidance. As shown in Table 2, our models are more capable of generating valid SMILES than other baseline models. Especially Top5-10 reactants, which are far ahead of other models. This indicates that incorporating 3D conformer information can assist the model in grasping the structures involved in chemical reactions, thus preventing the generation of SMILES that contravene chemical principles.
  (a) (b) Top-k accuracy (%)
1 3 5 10
- - 50.9 71.2 78.1 83.7
$\checkmark$ - 51.4 70.5 76.5 81.9
- $\checkmark$ 52.8 72.1 78.2 84.0
$\checkmark$ $\checkmark$ 53.6 73.7 80.2 86.1
Table 3: Ablation study of our method with reaction class unknown and random data augmentation. Module (a) is Atom-align Fusion; Module (b) is Distance-weighted Attention;
@@ 5.2 Ablation Study
To explore the impact of two modules, we design an ablation study under the settings of unknown reaction class and random SMILES data augmentation. The results are shown in Table 3 in the form of top-k accuracy and demonstrate necessity of each module for our method. The improvement in top-1 accuracy from ② to ① suggests that the model is able to use 3D positional information to infer the correct results with higher confidence after introducing module (a) Atom-align Fusion. Notably, the declines in top 3-10 accuracy is speculated on the reliance upon fixed 3D position embedding, which constrains the search space of model. Comparing ③ and ①, it can be seen that employing (b) Distance-weighted Attention improves performance in all top-k accuracy, even for top 3-10. We conjecture that redistributing attention enables the model to focus on internal structure of molecules, while Gaussian Kernel Function grants the model a larger search space. The results of ④ show that the combination of (a) and (b) has achieved the best performance of model, enabling it to focus on internal spatial structures while maintaining an extensive search space based on comprehensive 3D positional information.
@@ 5.3 Case Study
To verify that the proposed method our method can handle molecules with more intricate spatial structures, we chose 4 types of products as representatives:
@@
Regarding the 4 types of products, we experiment on our method and a Transformer without 3D information and choose the Top1-3 predictions for analysis, as shown in Figure 2. The rows of each subfigure correspond to our method and Transformer, respectively. Ground Truth and invalid SMILES are both highlighted in bold red.
The results show that when dealing with molecules of intricate spatial structures, our method can infer reasonable reactants with high confidence (often at top1, top2), while Transformer without 3D information finds it challenging to ascertain accurate results. Particularly in the case of Complex products, three predictions of our method could deduce skeleton and achieve the ground truth at top-1, whereas Transformer could not get any valid SMILES at top1-3. This further confirms the importance of introducing 3D information when dealing with molecules of intricate spatial structures.

### v2

@@ 4.3 SMILES Alignment
RetroSim Coley et al. (2017) 52.9 73.8 81.2 88.1 37.3 54.7 63.3 74.1
GLN Dai et al. (2019) 64.2 79.1 85.2 90.0 52.5 69.0 75.6 83.7
LocalRetro Chen and Jung (2021) 63.9 86.8 92.4 96.3 53.4 77.5 85.9 92.4
@@
RetroXpert Yan et al. (2020) 62.1 75.8 78.5 80.9 50.4 61.1 62.3 63.4
G2G Shi et al. (2020) 61.0 81.3 86.0 88.7 48.9 67.6 72.5 75.5
GraphRetro Somnath et al. (2020) 63.9 81.5 85.2 88.1 53.7 68.3 72.2 75.5
RetroPrime Wang et al. (2021) 64.8 81.6 85.0 86.9 51.4 70.8 74.0 76.1
MEGAN Sacha et al. (2021) 60.7 82.0 87.5 91.6 48.1 70.7 78.4 86.1
Graph2Edits Zhong et al. (2023) 67.1 87.5 91.5 93.6 55.1 77.3 83.4 89.4
@@
SCROP Zheng et al. (2019) 59.0 74.8 78.1 81.1 43.7 60.0 65.2 68.7
Aug. Transformer Tetko et al. (2020) - - - - 48.3 - 73.4 77.4
GTA Seo et al. (2021) - - - - 51.1 67.6 74.8 81.6
Graph2SMILES Tu and Coley (2022) - - - - 52.9 66.5 70.0 72.9
Retroformer Wan et al. (2022) 64.0 82.5 86.7 90.2 53.2 71.1 76.6 82.1
NAG2G Yao et al. (2024) 67.2 86.4 90.5 93.8 55.1 76.9 83.4 89.9
Retro3D (Ours) 67.6 87.2 91.4 94.5 55.5 77.2 83.4 89.1
Table 1: Top-k accuracy for Retrosynthesis prediction on USPTO-50K. The metric values of other models in the table are taken from their original papers. Best performance is in bold.
@@ 4.4 Overall Loss
The overall loss is composed of the language modeling R-Drop lossWu et al. (2021) $\mathcal{L}_{LM}$ for Retrosynthesis, along with the SMILES Alignment guidance loss $\mathcal{L}_{SA}$.
@@
$\mathcal{L}=\mathcal{L}_{LM}^{(CE)}+\alpha\mathcal{L}_{LM}^{(KL)}+\beta\mathcal{L}_{SA},$ (10)
where $\mathcal{L}_{LM}^{(CE)}$ is Cross-Entropy loss to minimize prediction error against actual outputs, $\mathcal{L}_{LM}^{(KL)}$ is KL-divergence in R-Drop loss to strive for outputs of models with varying Dropout to be as similar as possible. $\mathcal{L}_{SA}$ is Cross-Entropy loss between cross-attention score of last decoder layer and SAM that we mention at Section 4.3. We set $\alpha$ to 0.5, $\beta$ to 1.0.
Model Top-k Validity (%) Top-k Round-Trip (%)
1 3 5 10 1 3 5 10
Graph2SMILES 99.4 90.9 84.9 74.9 76.7 56.0 46.4 34.9
RetroPrime 98.9 98.2 97.1 92.5 79.6 59.6 50.3 40.4
Retroformer 99.2 98.5 97.4 96.7 78.9 72.0 67.1 57.2
NAG2G 99.7 98.6 97.1 92.9 - - - -
Retro3D (Ours) 99.8 99.6 99.1 97.0 90.7 81.3 75.7 69.0
Table 2: Top-k SMILES Validity / Round-Trip Accuracy for Retrosynthesis prediction on USPTO-50K with reaction class unknown.
(a) (b) (c) Top-k Accuracy (%) Top-k Validity (%)
1 3 5 10 1 3 5 10
- - - 50.2 70.6 76.0 80.1 99.1 98.4 97.7 95.6
- - $\checkmark$ 52.8 73.8 80.7 86.4 99.3 98.9 98.5 95.3
$\checkmark$ - $\checkmark$ 53.3 73.1 79.1 84.6 99.6 99.2 98.4 94.9
- $\checkmark$ $\checkmark$ 54.1 73.2 79.1 85.2 99.7 99.4 98.6 94.2
$\checkmark$ $\checkmark$ - 54.1 73.0 80.0 86.5 99.7 99.6 98.9 95.9
$\checkmark$ $\checkmark$ $\checkmark$ 55.5 77.2 83.4 89.1 99.8 99.6 99.1 97.0
Table 3: Ablation study of Retro3D with reaction class unknown on USPTO-50K dataset. Module (a) is Atom-align Fusion; Module (b) is Distance-weighted Attention and Module (c) is SMILES Alignment;
@@ 5 Experiments
We use the popular Retrosynthesis benchmark dataset USPTO-50K [^1] USPTO-50K and USPTO-FULL Dataset can be obtained in https://github.com/Hanjun-Dai/GLN Schneider et al. (2016), which contains 50,016 atom-mapped reactions, to evaluate our model. The same data split as Coley et al. (2017) is applied to our experiments, resulting in 40,008, 5,001, and 5,007 reactions for the training, validation, and test sets, respectively. Conformers are generated using RDKit with the stochastic optimization algorithm Merck Molecular Force Field (MMFF). We then apply the same algorithm in Wan et al. (2022) to extract the ground truth SMILES Alignment Map. We also conduct experiments on the larger dataset USPTO-FULL, details can be seen in Appendix B.
Setting.
@@
Our model is built on vanilla Transformer, consisting of 6 encoder layers and 6 decoder layers with 8 attention heads. The model dimension is set to 512. We employ the Adam optimizer Kingma and Ba (2014) with $(\beta_{1},\beta_{2})=(0.9,0.98)$ and train the model with batch size of 16. During training, we set an early-stop of 7 epochs under a maximum of 1000 epochs to obtain the 7 best models and average their parameters. Finally, the whole training process took approximately 30 hours on a single NVIDIA GeForce RTX 4090 GPU. Our code is available at https://anonymous.4open.science/r/Retro3D.
Evaluation.
@@
During inference, we use beam search Tillmann and Ney (2003) decoding strategy with a beam size is 10. After that, we evaluate our model with Top-k Accuracy, which considers it correct only if Top-k predicted reactants of the beam search results are the same as those in the original test set. We verify Top-k Validity based on whether RDKit can recognize the Top-k predicted reactants. Top-k Round-Trip Accuracy which measures the percentage of predicted reactants that can lead back to the original product.
Baseline.
@@
We compare our proposed method, Retro3D, with several strong baseline models from three types of retrosynthesis methods. We take RetroSim Coley et al. (2017), GLN Dai et al. (2019) and LocalRetro Chen and Jung (2021) to represent template-based methods. We take RetroXpert Yan et al. (2020), G2Gs Shi et al. (2020), GraphRetro Somnath et al. (2020), RetroPrime Wang et al. (2021), MEGAN Sacha et al. (2021) and Graph2Edits Zhong et al. (2023) to represent semi-template method. We take SCROP Zheng et al. (2019), Aug.Transformer Tetko et al. (2020), Graph2SMILES Zhong et al. (2023), GTA Seo et al. (2021) and Retroformer Wan et al. (2022), NAG2G Yao et al. (2024) to represent template-free method. We did not choose pre-trained methods for comparison as they are trained on larger datasets. We apply Root-align data augmentation Zhong et al. (2022) and conducted experiments on the proposed model Retro3D, choosing on-the-fly augmentation instead of directly expanding the training dataset.
5.1 Performance
Top-k Accuracy.
As shown in Table 1, when the reaction class is known, Retro3D achieves Top1-10 results that outperform all template-free/semi-template baselines and most state-of-the-art template-based baselines. Notably, our model exceeds both template-based and semi-template methods, despite their reliance on predefined templates or RDKit editing. When the reaction class is unknown, our model still achieves improvements in Top1-3 accuracy within the template-free category. Achieving high-confidence predictions (i.e., Top-1 acc), which narrows the search space for synthesis routes and enhances efficiency, makes our model particularly well-suited for real-world scenarios. Performance across each reaction classes in USPTO-50K dataset are provided in Appendix A. Result on USPTO-FULL dataset can be seen in Appendix B.
@@ 5.1 Performance
Figure 2: Case study for intricate structures. Each subfigure shows predictions of different methods: Retro3D (Top), Transformer (Bottom).
Top-k Validity.
We take three typical template-free methods Graph2SMILES, RetroPrime, Retroformer and NAG2G as baseline models to verify the predicted SMILES validity of our model. We refrained from selecting template-based and semi-template methods, as SMILES constructed based on template information ensure validity. In contrast, template-free methods often struggle to generate valid SMILES without extra chemical guidance. As shown in Table 2, our models are more capable of generating valid SMILES than other baseline models. Especially for Top5-10 reactants, which are far ahead of other models. This indicates that incorporating 3D conformer information can assist the model in grasping the structures involved in chemical reactions, thus preventing the generation of SMILES that violate chemical principles.
Top-k Round-Trip Accuracy.
To assess the accuracy of our predicted synthesis plans, we use the Molecule Transformer Schwaller et al. (2019) as the benchmark reaction prediction model and calculate the Top-k Round-Trip Accuracy. The results in Table 2 clearly show that our model significantly outperforms other baselines, achieving a remarkable 12% improvement in Top-1 Round-Trip Accuracy compared to the best baseline.
@@ 5.2 Ablation Study
To explore the impact of three modules, we design an ablation study under the settings of unknown reaction class. The results are shown in Table 3 in the form of top-k accuracy and demonstrate necessity of each module for Retro3D. ② highlights the necessity of module (c) SMILES Alignment. The improvement in top-1 accuracy from ③ to ② suggests that the model is capable of using 3D positional information to infer the correct results with higher confidence after introducing module (a) Atom-align Fusion. Notably, the declines in top3-10 accuracy is speculated on the reliance upon fixed 3D position embedding, which constrains the search space of model. Comparing ④ and ②, it can be seen that employing (b) Distance-weighted Attention improves performance in all top-k accuracy, even for top3-10. We conjecture that redistributing attention enables the model to focus on internal structure of molecules, while Gaussian Kernel Function grants the model a larger search space. The results of ⑥ show that the combination of (a), (b) and (c) has achieved the best performance of model, enabling it to focus on internal spatial structures while maintaining an extensive search space based on comprehensive 3D positional information.
@@ 5.3 Case Study
To verify that the proposed method Retro3D can handle molecules with more intricate spatial structures, we chose 4 types of products as representatives:
@@
We conducted experiments comparing Retro3D with a Vanilla Transformer, analyzing Top1-3 predictions, as illustrated in Figure 2. The rows in each subfigure represent the results of Retro3D and Transformer, respectively, with Ground Truth and invalid SMILES highlighted in bold red.
The results show that when dealing with molecules of intricate spatial structures, Retro3D can infer reasonable reactants with high confidence (often at top1, top2), while Transformer without 3D information struggles to generate accurate predictions. Notably, for Complex products, Retro3D successfully deduces the molecular skeleton and achieves the Ground Truth at Top-1, while the Vanilla Transformer fails to produce any valid SMILES within the Top-3 predictions.